OpenStudy (anonymous):

Which shows 5 6/9 as a simplified improper fraction? 17/3 45/9 51/9 5/3

OpenStudy (misty1212):

HI again!!

OpenStudy (misty1212):

is this \[5\tfrac{9}{9}\]?

OpenStudy (anonymous):

hi again:):)

OpenStudy (misty1212):

oops i mean \[\large 5\tfrac{6}{9}\]

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok then this is the same as \[5\tfrac{2}{3}\] when you reduce it

OpenStudy (misty1212):

then as a fraction \[\frac{3\times 5+2}{3}\]

OpenStudy (anonymous):

im confused

OpenStudy (misty1212):

oooh sorry do you know how to turn a mixed number in to an improper fraction? like \[2\tfrac{1}{2}=\frac{5}{2}\]?

OpenStudy (anonymous):

no i didnt learn it properly and i forgot how to do it

OpenStudy (misty1212):

ok so in our heads for \[2\tfrac{1}{2}\] you will have the two in the denomiantor, the numerator is two times two plus one or five

OpenStudy (misty1212):

for \[5\tfrac{2}{3}\] the denominator will be 3, the numerator will be three times five plus two three times five is fifteen, fifteen plus two is seventeen \[5\tfrac{2}{3}=\frac{17}{3}\]

OpenStudy (anonymous):

oh so you mulltiply your numerator with the number on the left

OpenStudy (anonymous):

ok i get it now thank u

OpenStudy (misty1212):

ok good \(\color\magenta\heartsuit\)